OpenStudy (anonymous):

2(x + 6) − 10 = 12 + 4(x − 1) Can someone explain to me how to do this equation

OpenStudy (andriod09):

2(x+6)-10=12+4(x-1) First, lets take the forst half of the equation. 2(x+6)-10.

OpenStudy (andriod09):

2(x+6)-10 [(2*x)+(2*6)]-10 2x+12-10.

OpenStudy (andriod09):

then: 12+4(x-1) 12+[(4*x)+((4*(-1))] 12+4x+-4

OpenStudy (andriod09):

then we have: 2x+12-10=12+4x+(-4) -2x = -2x

OpenStudy (andriod09):

12-10= 12+2x+(-4) -12 =-12 -10=2x+(-4)

OpenStudy (andriod09):

-10=2x+(-4) +04= +4 -06=2x -03=x

OpenStudy (andriod09):

thatas how you answer those kind of questions.